OpenStudy (anonymous):

graph -2x+4y=4 for the domain D: (-8,-4,0,4,8)

OpenStudy (anonymous):

@pooja195 @satellite73 @MathMath101 @peachpi

OpenStudy (anonymous):

@ganeshie8 @Hero

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

well we need to find first what the corresponding y values are for the listed x values... so if x=-8 then y=? use the equation -2x+4y=4 then you will ask yourself if x=-4 then y=? and so on... once we get all the points we just plot then done

OpenStudy (anonymous):

so we finding y now

OpenStudy (freckles):

have you found y for when x=-8 yet?

OpenStudy (anonymous):

y equals 2

OpenStudy (freckles):

\[-2x+4y=4 \\ \text{ so we have } x=-8 \\ -2(-8)+4y=4 \\ 16+4y=4 \\ 4y=4-16 \\ 4y=-12 \\ y=\frac{-12}{4}=-3\] are you sure?

OpenStudy (freckles):

because when I plug in -8 for x I get -3 for y

OpenStudy (anonymous):

o sorry

OpenStudy (freckles):

so we will plot (-8,-3) on our little graph paper now you go to the next number in the domain and plug in -4 for x to find what y is when x is -4

OpenStudy (anonymous):

1

OpenStudy (freckles):

close I got y=-1 \[-2(-4)+4y=4 \\ 8+4y=4 \\ 4y=4-8 \\ 4y=-4 \\ y=-1\]

OpenStudy (freckles):

so we will plot (-4,-1) on our graph also by the way instead of solving for y each time we can do it one time and just plug into that equation instead: \[-2x+4y=4 \\ \text{ add } 2x \text{ on both sides } 4y=2x+4 \\ \text{ divide both sides by 4} \\ y=\frac{2x+4}{4} \\ \text{ divide \top and bottom by 2 } \\ y=\frac{x+2}{2}\]

OpenStudy (freckles):

ok plug in the next x value

OpenStudy (freckles):

which is 0

OpenStudy (anonymous):

1

OpenStudy (freckles):

right so we will also plot (0,1)

OpenStudy (freckles):

just two more number to plug in

OpenStudy (freckles):

so much fun :p

OpenStudy (freckles):

just to summary what we have so far: \[(-8,-3) \\ (-4,-1) \\ (0,1) \\ (4,?) \\ (8,?)\]

OpenStudy (freckles):

\[(-8,-3) \\ (-4,-1) \\ (0,1) \\ (4,?) \\ (8,?) \\ y=\frac{x+2}{2} \\ (-8,-3) \\ (-4,-1) \\ (0,1) \\ (4,\frac{4+2}{2}) \\ (8, \frac{8+2}{2})\] simplify those bottom two

OpenStudy (freckles):

then plot all 5 coordinates and you are done

OpenStudy (anonymous):

-4

OpenStudy (freckles):

(4+2)/2=?

OpenStudy (freckles):

and (8+2)/2=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

5

OpenStudy (freckles):

right! \[(-8,-3) \\ (-4,-1) \\ (0,1) \\ (4,?) \\ (8,?) \\ y=\frac{x+2}{2} \\ (-8,-3) \\ (-4,-1) \\ (0,1) \\ (4,\frac{4+2}{2}) =(4,3)\\ (8, \frac{8+2}{2}) =(8,5)\]

OpenStudy (freckles):

can you plot the points: (-8,-3) (-4,-1) (4,3) (8,5)

OpenStudy (freckles):

oops and (0,1)

OpenStudy (freckles):

I will plot one point if you are unsure how to plot

OpenStudy (freckles):

(-8,-3) look at the first number which is -8 since the number is negative that means we will go left of the origin (the center aka (0,0) ) we will count 8 units to the left of the origin ... the the number which is -3 since the number is negative that means we will go down from the origin we will count down 3 units from the origin the type graphing we are doing is called rectangular ...